Show that the conicoid x^2+2y^2+2yz+2x+4y+8z+1=0 is central. Find the new equation of the conicoid if the origin is shifted to its centre.
1
Expert's answer
2020-12-07T10:53:27-0500

A conicoid given by equation "ax^2+by^2+cz^2+2fyz+2gxz+2hxy+2ux+2vy+2wz+d=0"

is central if the system


"\begin{cases}

ax+hy+gz+u=0 \\

hx+by+fz+v=0 \\

gx+fy+cz+w=0

\end{cases}"


has a unique solution "(x_0,y_0,z_0)". In this case the point "(x_0,y_0,z_0)" is a center of the conicoid.


In our case, for a conicoid "x^2+2y^2+2yz+2x+4y+8z+1=0" the system is the following:


"\begin{cases}
x+1=0 \\
2y+z+2=0 \\
y+4=0
\end{cases}"


It is equivalent to


"\begin{cases}
x=-1 \\
y=-4 \\
z=6
\end{cases}"


Therefore, the conicoid is central with center at "O'(-1,-4,6)".


Let us find the new equation of the conicoid if the origin is shifted to its centre.


"\begin{cases}
x=x'-1 \\
y=y'-4 \\
z=z'+6
\end{cases}"


"(x'-1)^2+2(y'-4)^2+2(y'-4)(z'+6)+2(x'-1)+4(y'-4)+8(z'+6)+1=0"


"(x')^2-2x'+1+2(y')^2-16y'+32+2y'z'+12y'-8z'-48+2x'-2+4y'-16+8z'+48+1=0"


Therefore, the new equation of the conicoid if the origin is shifted to its centre is the following:


"(x')^2+2(y')^2+2y'z'+16=0"